A cyclist overcomes a resistive force or 30 N in order to cycle 30 m. It takes her 6 seconds to cycle this distance. Calculate the power of the cyclist


1
Expert's answer
2021-01-19T20:01:09-0500

By the definition of the power, we get:


"P=\dfrac{W}{t}=\dfrac{Fd}{t}=\dfrac{30\ N\cdot 30\ m}{6\ s}=150\ W."

Answer:

"P=150\ W."